mathslover (mathslover):

Logarithm Question (4) Solve : \(5.x^{\log_7 2} + 2^{\log_7 x} = 24\)

OpenStudy (shadowlegendx):

Calculus go away

mathslover (mathslover):

This time, I have no clue for the first step.

Miracrown (miracrown):

is that 5.x or is the 5 being multiplied?

mathslover (mathslover):

5 . x = 5 * x = 5(x) ...

OpenStudy (shadowlegendx):

Well doesnt log stand for something? o-o

Miracrown (miracrown):

 so the 5 is outside of the x so it is like that would you agree?